OpenStudy (anonymous):

How would u find the equation for this hyperbola- 4x^2-y^2=0 ????

OpenStudy (oaktree):

That's not a hyperbola. It's a normal function. Add y^2: y^2=4x^2. Square root, and y=2x.

OpenStudy (amistre64):

+- ...

OpenStudy (amistre64):

its a degenerate hyperbola

OpenStudy (anonymous):

but i'm trying to make a hyperbola out of it..

OpenStudy (anonymous):

who you calling a degenerate?

OpenStudy (amistre64):

ellen lol

OpenStudy (anonymous):

so how would i do it...

OpenStudy (anonymous):

ho ho ho

OpenStudy (amistre64):

y = +- 2x

OpenStudy (amistre64):

a degenerate hyperbola is an X shape

OpenStudy (anonymous):

\[4x^2=y^2\] \[y=2x \text{ or } y = -2x\] a nice pair of lines, what amistre said degeneration X

OpenStudy (anonymous):

ok. that makes sense buthow would i graph it...

OpenStudy (amistre64):

with 2 lines; y = 2x and y=-2x

OpenStudy (anonymous):

k and the center would be the origin?

OpenStudy (amistre64):

well, both y intercepts amount to 0, so yes

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%5E2%3D4x%5E2

OpenStudy (anonymous):

cool. and would there be a foci and/or vertices?

OpenStudy (amistre64):

the verts touch and are at the same points; which is why this thing is degenerate; youd have to play about for the focuses tho; i cant be sure of them unless they are at the center as well

OpenStudy (anonymous):

k thank u sooooooo much! ur awesome!

OpenStudy (amistre64):

im awesome AND i smell like irish spring lol

OpenStudy (anonymous):

haha random but awesome :)